OpenStudy (anonymous):

Graph y = {3 if x is less than or equal to -1} {2x if -1 is less than x is less than 5} {(3-x) if x is greater than or equal to 5}

OpenStudy (aum):



OpenStudy (anonymous):

thank you, but can you also please explain to me how you got that?

OpenStudy (aum):

There are 3 x-intervals and the function has different equations in each of the interval. When \(x \le -1\), y = 3 y = 3 is a horizontal line where the y-value remains the same no matter what the x-value is. But y = 3 only when \(x \le -1\). So we draw a horizontal line at y = 3 and stop it when it reaches x = -1. Also, since x = -1 is included in \(x \le -1\), we put a dark circle at the end to indicate that the point is included in the line. Is this part clear?

OpenStudy (aum):



OpenStudy (anonymous):

Yes, thank you

OpenStudy (aum):

Are the other two parts of the graph clear too?